1. In a reaction, 5.3 g of sodium carbonate reacted with 6 g of ethanoic acid. The products were 2.2 g of carbon dioxide, 0.9 g water and 8.2 g of sodium ethanoate. Show that these observations are in agreement with the law of conservation of mass.
Sodium carbonate + ethanoic acid " sodium ethanoate + carbon dioxide + water

Grade:12th pass

1 Answers

Pawan Prajapati
3 years ago
In the given reaction, sodium carbonate reacts with ethanoic acid to produce sodium ethanoate, carbon dioxide, and water. Sodium carbonate + Ethanoic acid " Sodium ethanoate + Carbon dioxide + Water Mass of sodium carbonate = 5.3 g (Given) Mass of ethanoic acid = 6 g (Given) Mass of sodium ethanoate = 8.2 g (Given) Mass of carbon dioxide = 2.2 g (Given) Mass of water = 0.9 g (Given) Now, total mass before the reaction =(5.3 + 6) g = 11.3 g And, total mass after the reaction = (8.2 + 2.2 + 0.9) g = 11.3 g Total mass before the reaction = Total mass after the reaction Hence, the given observations are in agreement with the law of conservation of mass.